Jefferson Lab HALOG ARCHIVES Logbook for March 00 - All EXCEPT Automatic Entries

  • Main INDEX

    Friday

  • 37515 03/31/00 23:37 All Kin#102 Run Summary.
  • 37514 03/31/00 23:23 kuss E89044 acknowledgements II
  • 37513 03/31/00 22:31 Quemener & K E89044 finished
  • 37512 03/31/00 22:05 Marcucci shift summary
  • 37511 03/31/00 21:05 adev Accepted_Events_Run_3100
  • 37509 03/31/00 20:13 adev Accepted_Events_Run_3099
  • 37506 03/31/00 19:47 adaq beam strip charts
  • 37505 03/31/00 19:45 adaq magnet strip charts
  • 37504 03/31/00 18:48 adev Accepted_Events_Run_3098
  • 37500 03/31/00 17:22 adev Accepted_Events_Run_3097
  • 37498 03/31/00 17:16 adev Accepted_Events_Run_3096
  • 37497 03/31/00 17:16 adev Accepted_Events_Run_3096
  • 37492 03/31/00 16:08 adev Accepted_Events_Run_3095
  • 37489 03/31/00 16:01 F.Benmokhtar Shift Summary
  • 37487 03/31/00 14:52 F.Benmokhtar Beam Stripcharts
  • 37486 03/31/00 14:46 adev Accepted_Events_Run_3094
  • 37483 03/31/00 14:45 adev Accepted_Events_Run_3093
  • 37481 03/31/00 14:39 F.Benmokhtar Magnet Screen (Probe#3)
  • 37479 03/31/00 13:57 adev Accepted_Events_Run_3092
  • 37476 03/31/00 13:29 F.Benmokhtar Magnet Screen for kin#102
  • 37475 03/31/00 12:00 adev Accepted_Events_Run_3090
  • 37474 03/31/00 11:57 Marat&Fatiha Delta Scan runs
  • 37472 03/31/00 11:48 F.Benmokhtar Beam Stripcharts
  • 37470 03/31/00 11:30 adev Accepted_Events_Run_3089
  • 37467 03/31/00 11:26 adev Accepted_Events_Run_3088
  • 37463 03/31/00 11:02 adev Accepted_Events_Run_3087
  • 37461 03/31/00 11:00 adev Accepted_Events_Run_3086
  • 37458 03/31/00 10:36 F.Benmokhtar Harp scan Run#763
  • 37456 03/31/00 08:12 adev Accepted_Events_Run_3085
  • 37455 03/31/00 08:07 R. Michaels runstart info wrong - run 3085
  • 37453 03/31/00 08:07 M. Rvachev Shift summary
  • 37451 03/31/00 07:58 adev Accepted_Events_Run_3084
  • 37448 03/31/00 07:55 adev Accepted_Events_Run_3083
  • 37447 03/31/00 07:54 adev Accepted_Events_Run_3082
  • 37442 03/31/00 07:33 adev Accepted_Events_Run_3081
  • 37440 03/31/00 07:20 R. Michaels deadtime correlated to ESPACE
  • 37438 03/31/00 07:06 adev Accepted_Events_Run_3080
  • 37436 03/31/00 06:54 adev Accepted_Events_Run_3079
  • 37432 03/31/00 06:29 adev Accepted_Events_Run_3078
  • 37429 03/31/00 05:29 adev Accepted_Events_Run_3077
  • 37427 03/31/00 05:07 adev Accepted_Events_Run_3076
  • 37423 03/31/00 04:49 adev Accepted_Events_Run_3075
  • 37422 03/31/00 04:47 M. Rvachev Magnet stripcharts
  • 37419 03/31/00 04:36 adev Accepted_Events_Run_3074
  • 37417 03/31/00 04:36 M. Rvachev Magnets screen
  • 37415 03/31/00 04:33 M. Rvachev Beam stripcharts
  • 37414 03/31/00 04:25 M. Rvachev OTR check
  • 37413 03/31/00 04:19 adev Accepted_Events_Run_3073
  • 37410 03/31/00 04:05 adev Accepted_Events_Run_3072
  • 37406 03/31/00 03:49 adev Accepted_Events_Run_3071
  • 37405 03/31/00 03:42 adev Accepted_Events_Run_3070
  • 37401 03/31/00 03:29 adev Accepted_Events_Run_3069
  • 37397 03/31/00 03:22 adev Accepted_Events_Run_3068
  • 37395 03/31/00 03:11 adev Accepted_Events_Run_3067
  • 37392 03/31/00 03:04 adev Accepted_Events_Run_3066
  • 37389 03/31/00 02:55 adev Accepted_Events_Run_3065
  • 37386 03/31/00 02:37 adev Accepted_Events_Run_3064
  • 37384 03/31/00 02:34 R. Michaels estimates of trigger efficiency
  • 37383 03/31/00 02:31 adev Accepted_Events_Run_3063
  • 37381 03/31/00 02:28 R. Michaels run 3063 - current drift
  • 37379 03/31/00 02:15 adev Accepted_Events_Run_3061
  • 37373 03/31/00 01:47 adev Accepted_Events_Run_3060
  • 37371 03/31/00 01:29 adev Accepted_Events_Run_3059
  • 37368 03/31/00 01:06 adev Accepted_Events_Run_3058
  • 37366 03/31/00 00:56 R. Michaels note re: E-arm momentum
  • 37364 03/31/00 00:44 adev Accepted_Events_Run_3057
  • 37362 03/31/00 00:40 adev Accepted_Events_Run_3055
  • 37360 03/31/00 00:36 adev Accepted_Events_Run_3056
  • 37356 03/31/00 00:23 kuss Shift summary
  • 37355 03/31/00 00:10 adev Accepted_Events_Run_3054
  • 37350 03/31/00 00:01 adev Accepted_Events_Run_3053
  • Thursday

  • 37348 03/30/00 23:43 adev Accepted_Events_Run_3052
  • 37345 03/30/00 23:40 adev Accepted_Events_Run_3051
  • 37342 03/30/00 23:23 adev Accepted_Events_Run_3050
  • 37338 03/30/00 23:05 adev Accepted_Events_Run_3049
  • 37336 03/30/00 22:47 adev Accepted_Events_Run_3047
  • 37335 03/30/00 22:47 adev Accepted_Events_Run_3048
  • 37328 03/30/00 22:28 adev Accepted_Events_Run_3046
  • 37326 03/30/00 22:15 adev Accepted_Events_Run_3045
  • 37324 03/30/00 22:00 adev Accepted_Events_Run_3044
  • 37320 03/30/00 21:28 adev Accepted_Events_Run_3043
  • 37318 03/30/00 21:18 adev Accepted_Events_Run_3042
  • 37314 03/30/00 20:43 adev Accepted_Events_Run_3041
  • 37312 03/30/00 20:30 fissum magnet stripcharts
  • 37311 03/30/00 20:29 fissum beam stripcharts
  • 37310 03/30/00 20:26 fissum magnet stripcharts
  • 37308 03/30/00 20:03 fissum otr results
  • 37307 03/30/00 19:51 adev Accepted_Events_Run_3040
  • 37305 03/30/00 19:43 adev Accepted_Events_Run_3039
  • 37301 03/30/00 19:23 fissum harp scan results (761)
  • 37299 03/30/00 19:11 adev Accepted_Events_Run_3038
  • 37295 03/30/00 17:38 adev Accepted_Events_Run_3037
  • 37293 03/30/00 13:56 adev Accepted_Events_Run_3036
  • 37290 03/30/00 08:00 M. Rvachev Shift summary
  • 37288 03/30/00 07:14 M. Rvachev List of 1254 MeV elastic runs (kin 91)
  • 37287 03/30/00 07:06 adev Accepted_Events_Run_3035
  • 37285 03/30/00 06:54 adev Accepted_Events_Run_3034
  • 37284 03/30/00 06:46 R. Michaels ER problems at high rates
  • 37281 03/30/00 06:27 adev Accepted_Events_Run_3033
  • 37278 03/30/00 06:09 M. Rvachev Magnet stripcharts
  • 37277 03/30/00 06:07 M. Rvachev Beam stripcharts
  • 37275 03/30/00 05:56 adev Accepted_Events_Run_3032
  • 37272 03/30/00 05:42 adev Accepted_Events_Run_3031
  • 37269 03/30/00 05:30 M. Rvachev Moved to elastic scattering point; magnets screen
  • 37268 03/30/00 05:27 adev Accepted_Events_Run_3030
  • 37265 03/30/00 04:49 M. Rvachev List of kinematics 23 runs
  • 37264 03/30/00 04:45 adev Accepted_Events_Run_3029
  • 37261 03/30/00 04:26 adev Accepted_Events_Run_3028
  • 37258 03/30/00 04:17 adev Accepted_Events_Run_3027
  • 37255 03/30/00 04:00 adev Accepted_Events_Run_3026
  • 37253 03/30/00 03:48 M. Rvachev Magnets screen
  • 37251 03/30/00 03:13 adev Accepted_Events_Run_3024
  • 37248 03/30/00 02:17 adev Accepted_Events_Run_3023
  • 37245 03/30/00 01:15 adev Accepted_Events_Run_3022
  • 37242 03/30/00 00:15 adev Accepted_Events_Run_3021
  • Wednesday

  • 37238 03/29/00 23:37 adev Accepted_Events_Run_3020
  • 37235 03/29/00 23:09 adev Accepted_Events_Run_3019
  • 37232 03/29/00 22:31 adev Accepted_Events_Run_3018
  • 37229 03/29/00 22:02 bogdanw beam optics
  • 37228 03/29/00 22:00 bogdanw magnets stripcharts
  • 37227 03/29/00 21:52 adev Accepted_Events_Run_3017
  • 37222 03/29/00 21:07 adev Accepted_Events_Run_3016
  • 37219 03/29/00 19:40 adev Accepted_Events_Run_3015
  • 37217 03/29/00 19:11 adev Accepted_Events_Run_3014
  • 37215 03/29/00 18:45 adev Accepted_Events_Run_3013
  • 37210 03/29/00 18:04 adev Accepted_Events_Run_3012
  • 37208 03/29/00 17:29 adev Accepted_Events_Run_3011
  • 37206 03/29/00 17:28 bogdanw OTR
  • 37203 03/29/00 16:11 mitchell Day shift summary
  • 37202 03/29/00 16:02 F.Benmokhtar Beam Stripcharts.
  • 37201 03/29/00 16:00 F.Benmokhtar Harp scan
  • 37200 03/29/00 16:00 mitchell ep energy online result
  • 37199 03/29/00 15:49 adev Accepted_Events_Run_3010
  • 37196 03/29/00 15:03 adev Accepted_Events_Run_3009
  • 37193 03/29/00 14:31 adev Accepted_Events_Run_3008
  • 37190 03/29/00 14:17 adev Accepted_Events_Run_3007
  • 37187 03/29/00 14:15 adev Accepted_Events_Run_3006
  • 37184 03/29/00 12:36 F.Benmokhtar Magnet Stipcharts
  • 37183 03/29/00 12:35 F.Benmokhtar Beam Stipcharts
  • 37182 03/29/00 12:11 adev Accepted_Events_Run_3005
  • 37178 03/29/00 11:59 adev Accepted_Events_Run_3004
  • 37175 03/29/00 11:40 adev Accepted_Events_Run_3003
  • 37172 03/29/00 11:28 adev Accepted_Events_Run_3002
  • 37167 03/29/00 11:04 adev Accepted_Events_Run_3000
  • 37164 03/29/00 11:02 adev Accepted_Events_Run_2999
  • 37161 03/29/00 10:20 adev Accepted_Events_Run_2998
  • 37158 03/29/00 10:05 F.Benmokhtar Harp Scan raster off.
  • 37157 03/29/00 09:53 mitchell OTR after adjustments
  • 37156 03/29/00 09:34 mitchell OTR after restoration
  • 37154 03/29/00 09:31 adev Accepted_Events_Run_2997
  • 37151 03/29/00 09:28 adev Accepted_Events_Run_2996
  • 37149 03/29/00 08:03 Z. Chai Owl Shift Summary
  • 37147 03/29/00 06:16 Z. Chai magnets strip plot
  • Tuesday

  • 37145 03/28/00 23:58 Punjabi Shift Summary
  • 37144 03/28/00 21:21 adev Accepted_Events_Run_2995
  • 37142 03/28/00 21:10 Punjabi OTR scan
  • 37140 03/28/00 20:59 Perdrisat more strip charts
  • 37139 03/28/00 20:55 Perdrisat strip charts
  • 37138 03/28/00 20:54 adev Accepted_Events_Run_2994
  • 37136 03/28/00 20:50 Perdrisat harp scan
  • 37134 03/28/00 19:50 adev Accepted_Events_Run_2993
  • 37131 03/28/00 19:37 adev Accepted_Events_Run_2992
  • 37128 03/28/00 18:31 Higinbotham Arc Energy at 2 Pass Beam
  • 37127 03/28/00 16:02 jpchen Shift Summary
  • 37125 03/28/00 15:37 jpchen Back to beam permit
  • 37124 03/28/00 15:22 jpchen Beam trip due to PSS problem
  • 37123 03/28/00 13:04 jpchen jpchen
  • 37122 03/28/00 13:02 jpchen Arc scanner high voltage set
  • 37121 03/28/00 13:01 jpchen Bogie crate rebooted once
  • 37120 03/28/00 12:51 jpchen e-spectrometer is at 51.45 degree
  • 37119 03/28/00 11:52 jpchen e-Spectrometer motion problem update
  • 37118 03/28/00 09:36 jpchen e-spectrometer motion probelm
  • 37117 03/28/00 09:11 jpchen Magnets setting for e kin 23
  • 37116 03/28/00 08:42 jpchen Target in 12C position
  • 37115 03/28/00 08:33 jpchen Spectrometer strip charts
  • 37114 03/28/00 08:32 jpchen beam Energy strip chart
  • 37113 03/28/00 08:29 jpchen strip chart BPMs
  • 37112 03/28/00 08:28 jpchen Magnet Settings
  • 37111 03/28/00 08:10 adev Accepted_Events_Run_2991
  • 37110 03/28/00 08:09 M. Rvachev List of kinematics 32 runs
  • 37107 03/28/00 08:03 M. Rvachev Shift summary
  • 37105 03/28/00 07:04 adev Accepted_Events_Run_2990
  • 37103 03/28/00 06:24 adev Accepted_Events_Run_2989
  • 37100 03/28/00 05:17 M. Rvachev Magnets screen
  • 37099 03/28/00 05:17 M. Rvachev Magnet stripcharts
  • 37098 03/28/00 05:16 M. Rvachev Beam stripcharts
  • 37096 03/28/00 05:08 adev Accepted_Events_Run_2988
  • 37093 03/28/00 03:42 adev Accepted_Events_Run_2987
  • 37089 03/28/00 03:08 adev Accepted_Events_Run_2986
  • 37088 03/28/00 02:55 adev Accepted_Events_Run_2985
  • 37084 03/28/00 01:20 adev Accepted_Events_Run_2984
  • 37082 03/28/00 00:04 adev Accepted_Events_Run_2983
  • 37081 03/28/00 00:01 Adaq shift summary
  • Monday

  • 37078 03/27/00 22:42 adev Accepted_Events_Run_2982
  • 37073 03/27/00 21:26 adev Accepted_Events_Run_2981
  • 37071 03/27/00 20:43 fissum update
  • 37070 03/27/00 20:07 adev Accepted_Events_Run_2980
  • 37067 03/27/00 18:46 adev Accepted_Events_Run_2979
  • 37064 03/27/00 18:09 fissum magnet stripcharts
  • 37063 03/27/00 18:08 fissum beam stripcharts
  • 37062 03/27/00 18:05 fissum otr 2979, see elog 105687, 105688
  • 37061 03/27/00 17:23 fissum otr during 2979
  • 37059 03/27/00 17:07 fissum ended run 2978
  • 37057 03/27/00 17:04 fissum 2976 - no endrun, 2977 - no data
  • 37056 03/27/00 17:03 adev Accepted_Events_Run_2978
  • 37055 03/27/00 17:01 fissum coda restart
  • 37053 03/27/00 16:24 adev Accepted_Events_Run_2975
  • 37050 03/27/00 16:13 fissum shift changeover
  • 37048 03/27/00 16:00 LeRose Shift Summary
  • 37047 03/27/00 15:06 adev Accepted_Events_Run_2974
  • 37044 03/27/00 14:28 LeRose More Current?
  • 37043 03/27/00 14:17 LeRose Magnet Strip Charts
  • 37042 03/27/00 14:16 LeRose Beam Strip Charts
  • 37040 03/27/00 13:39 adev Accepted_Events_Run_2973
  • 37037 03/27/00 12:40 adev Accepted_Events_Run_2972
  • 37035 03/27/00 11:28 LeRose Beam Strip Tool Restarted
  • 37034 03/27/00 11:23 LeRose Beam Back
  • 37032 03/27/00 10:45 LeRose Beam Still OFF
  • 37031 03/27/00 09:49 LeRose Beam Strip Charts
  • 37030 03/27/00 09:48 LeRose Magnet Strip Charts
  • 37029 03/27/00 09:15 LeRose Beam Off
  • 37028 03/27/00 09:08 adev Accepted_Events_Run_2971
  • 37025 03/27/00 08:03 M. Rvachev Shift summary
  • 37024 03/27/00 08:00 adev Accepted_Events_Run_2970
  • 37021 03/27/00 07:20 M. Rvachev Lost BPM signals on hac
  • 37019 03/27/00 06:36 adev Accepted_Events_Run_2969
  • 37017 03/27/00 06:35 adev Accepted_Events_Run_2968
  • 37014 03/27/00 05:37 F.Benmokhtar P0_e and P0_h stript charts
  • 37013 03/27/00 05:36 F.Benmokhtar Beam stript charts
  • 37012 03/27/00 05:24 adev Accepted_Events_Run_2967
  • 37009 03/27/00 04:13 adev Accepted_Events_Run_2966
  • 37005 03/27/00 02:52 M. Rvachev OTR check
  • 37004 03/27/00 02:49 adev Accepted_Events_Run_2965
  • 37002 03/27/00 01:37 adev Accepted_Events_Run_2964
  • Sunday

  • 36996 03/26/00 23:23 adev Accepted_Events_Run_2963
  • 36995 03/26/00 22:39 fissum magnet stripcharts
  • 36994 03/26/00 22:37 fissum beam stripcharts
  • 36991 03/26/00 22:22 adev Accepted_Events_Run_2962
  • 36990 03/26/00 22:20 adev Accepted_Events_Run_2961
  • 36985 03/26/00 21:04 adev Accepted_Events_Run_2960
  • 36982 03/26/00 19:49 adev Accepted_Events_Run_2959
  • 36981 03/26/00 19:41 Ulmer OTR - Run 2959
  • 36979 03/26/00 19:34 adev Accepted_Events_Run_2958
  • 36977 03/26/00 19:05 Ulmer magnet stripcharts
  • 36976 03/26/00 19:04 Ulmer Beam stripcharts
  • 36975 03/26/00 18:21 adev Accepted_Events_Run_2957
  • 36972 03/26/00 17:11 adev Accepted_Events_Run_2956
  • 36968 03/26/00 16:06 Quemener Shift Summary
  • 36967 03/26/00 15:49 adev Accepted_Events_Run_2955
  • 36963 03/26/00 14:32 adev Accepted_Events_Run_2954
  • 36959 03/26/00 12:37 adev Accepted_Events_Run_2952
  • 36955 03/26/00 10:37 adev Accepted_Events_Run_2950
  • 36953 03/26/00 10:27 adaq beam strip charts
  • 36952 03/26/00 10:27 adaq magnet strip charts
  • 36951 03/26/00 09:49 Punjabi otr from elog
  • 36949 03/26/00 09:02 adev Accepted_Events_Run_2949
  • 36945 03/26/00 08:51 adev Accepted_Events_Run_2948
  • 36943 03/26/00 08:00 M. Rvachev Shift summary
  • 36941 03/26/00 07:36 adev Accepted_Events_Run_2947
  • 36939 03/26/00 06:28 pibero Magnet StripCharts
  • 36937 03/26/00 06:16 M. Rvachev OTR check
  • 36936 03/26/00 06:13 pibero Beam strip charts
  • 36935 03/26/00 06:10 adev Accepted_Events_Run_2946
  • 36934 03/26/00 05:24 adev Accepted_Events_Run_2945
  • 36930 03/26/00 04:13 adev Accepted_Events_Run_2944
  • 36927 03/26/00 03:02 adev Accepted_Events_Run_2943
  • 36924 03/26/00 01:56 adev Accepted_Events_Run_2942
  • 36921 03/26/00 00:57 pibero magnet screen
  • 36919 03/26/00 00:27 adev Accepted_Events_Run_2941
  • 36915 03/26/00 00:18 adev Accepted_Events_Run_2940
  • 36913 03/26/00 00:06 adev Accepted_Events_Run_2939
  • Saturday

  • 36909 03/25/00 23:55 adev Accepted_Events_Run_2938
  • 36906 03/25/00 23:40 Ulmer Magnets screen
  • 36905 03/25/00 21:07 adev Accepted_Events_Run_2937
  • 36902 03/25/00 20:49 adev Accepted_Events_Run_2936
  • 36899 03/25/00 20:32 adev Accepted_Events_Run_2935
  • 36897 03/25/00 20:30 all Kin#31 summary
  • 36893 03/25/00 19:03 adev Accepted_Events_Run_2932
  • 36891 03/25/00 18:42 Ulmer Magnets screen
  • 36890 03/25/00 18:37 Ulmer 89044_hspect_p0
  • 36889 03/25/00 18:36 Ulmer 89044_espect_p0
  • 36888 03/25/00 18:35 Ulmer E89044_beam;_denergy
  • 36885 03/25/00 17:47 adev Accepted_Events_Run_2931
  • 36882 03/25/00 16:01 aniol shift summary
  • 36881 03/25/00 15:13 Aniol kin 31, 27.75hrs of data
  • 36879 03/25/00 09:02 adev Accepted_Events_Run_2930
  • 36877 03/25/00 08:26 adev Accepted_Events_Run_2929
  • 36875 03/25/00 08:02 M. Rvachev Shift summary
  • 36872 03/25/00 06:15 M. Rvachev Magnets screen
  • 36871 03/25/00 06:15 M. Rvachev Beam stripcharts
  • 36870 03/25/00 06:14 M. Rvachev Magnet stripcharts
  • 36869 03/25/00 04:12 adev Accepted_Events_Run_2928
  • 36865 03/25/00 03:23 adev Accepted_Events_Run_2927
  • 36864 03/25/00 03:22 adev Accepted_Events_Run_2926
  • 36861 03/25/00 02:05 adev Accepted_Events_Run_2925
  • 36858 03/25/00 00:50 adev Accepted_Events_Run_2924
  • Friday

  • 36854 03/24/00 23:35 adev Accepted_Events_Run_2923
  • 36851 03/24/00 22:19 adev Accepted_Events_Run_2922
  • 36848 03/24/00 21:22 Ulmer 3He(e,e'd) - Kinematics 31
  • 36846 03/24/00 20:58 Chang Magnet stript charts
  • 36845 03/24/00 20:56 Chang Beam optics stript charts
  • 36844 03/24/00 20:50 adev Accepted_Events_Run_2921
  • 36842 03/24/00 19:32 adev Accepted_Events_Run_2920
  • 36839 03/24/00 18:13 adev Accepted_Events_Run_2919
  • 36836 03/24/00 17:51 Chang Magnets
  • 36835 03/24/00 17:49 Chang Optics
  • 36834 03/24/00 17:47 Chang Beam energy strip charts
  • 36833 03/24/00 16:56 adev Accepted_Events_Run_2918
  • 36829 03/24/00 16:12 glashausser shift summary
  • 36827 03/24/00 14:20 glashausser spot++ events
  • 36825 03/24/00 13:57 adev Accepted_Events_Run_2916
  • 36821 03/24/00 13:54 adev Accepted_Events_Run_2915
  • 36818 03/24/00 09:21 krishni Magnet screens
  • 36817 03/24/00 09:21 krishniw Magnet screens
  • 36816 03/24/00 09:18 krishniw Beam strip charts
  • 36814 03/24/00 09:06 adev Accepted_Events_Run_2914
  • 36813 03/24/00 08:06 M. Rvachev Shift summary
  • 36810 03/24/00 07:58 adev Accepted_Events_Run_2913
  • 36809 03/24/00 07:02 M. Rvachev H-arm scintillator HV tripped off
  • 36808 03/24/00 06:36 adev Accepted_Events_Run_2912
  • 36805 03/24/00 05:15 M. Rvachev magnet stripcharts
  • 36804 03/24/00 05:08 m itchell strip charts
  • 36803 03/24/00 05:07 adev Accepted_Events_Run_2911
  • 36800 03/24/00 01:56 adev Accepted_Events_Run_2910
  • 36797 03/24/00 00:05 adev Accepted_Events_Run_2909
  • 36794 03/24/00 00:02 G. Quemener Shift summary
  • Thursday

  • 36792 03/23/00 23:01 Ulmer E-ARM gamma (probe 3)
  • 36790 03/23/00 22:22 adev Accepted_Events_Run_2908
  • 36788 03/23/00 21:15 Adaq end of run problem
  • 36787 03/23/00 21:06 adev Accepted_Events_Run_2907
  • 36783 03/23/00 19:36 adev Accepted_Events_Run_2906
  • 36781 03/23/00 19:32 adaq beam energy strip charts
  • 36780 03/23/00 19:29 adaq magnets strip charts
  • 36778 03/23/00 18:16 adev Accepted_Events_Run_2905
  • 36776 03/23/00 16:58 adev Accepted_Events_Run_2904
  • 36772 03/23/00 16:00 Kelly shift summary
  • 36771 03/23/00 15:13 eric Magnet strip chart
  • 36769 03/23/00 14:46 F.Benmokhtar Harp scan Scanner#5 results.
  • 36768 03/23/00 14:43 F.Benmokhtar Arc Scan Run#706, Harp Scan Run#707
  • 36767 03/23/00 13:25 adev Accepted_Events_Run_2903
  • 36764 03/23/00 13:08 adev Accepted_Events_Run_2902
  • 36762 03/23/00 12:56 Kelly OTR
  • 36761 03/23/00 12:49 F,Benmokhtar Beam and Energy Stript charts.
  • 36760 03/23/00 12:46 F.Benmokhtar Harp Scan raster off Run#705
  • 36759 03/23/00 12:41 F.Benmokhtar Spot size
  • 36757 03/23/00 11:45 F.Benmokhtar Target IOC Rebooted
  • 36756 03/23/00 10:18 F.Benmokhtar E_P0 and H_P0 Stript charts.
  • 36755 03/23/00 10:17 F. Benmokhta Energy and Beam Stript charts.
  • 36754 03/23/00 10:13 F.Benmokhtar Magnet setting for kin#31
  • 36753 03/23/00 09:02 adev Accepted_Events_Run_2901
  • 36750 03/23/00 08:14 Aniol shift summary
  • 36749 03/23/00 08:09 adev Accepted_Events_Run_2900
  • 36746 03/23/00 08:00 aniol Kinematics 12 run summary
  • 36745 03/23/00 05:50 Aniol kill coda
  • 36742 03/23/00 04:13 adev Accepted_Events_Run_2897
  • 36739 03/23/00 04:08 Aniol magnet strip charts
  • 36738 03/23/00 04:07 Aniol beam strip charts
  • 36736 03/23/00 01:24 adev Accepted_Events_Run_2896
  • 36733 03/23/00 00:56 adev Accepted_Events_Run_2895
  • 36731 03/23/00 00:05 LeRose Swing Shift Summary
  • Wednesday

  • 36729 03/22/00 23:13 LeRose Beam Back
  • 36727 03/22/00 22:52 adev Accepted_Events_Run_2894
  • 36725 03/22/00 22:44 LeRose Pulsed Beam
  • 36724 03/22/00 21:47 LeRose More Strip Charts
  • 36723 03/22/00 21:46 LeRose Strip Charts
  • 36722 03/22/00 21:32 LeRose Beam is back
  • 36721 03/22/00 21:14 LeRose Beam OFF
  • 36720 03/22/00 21:14 adev Accepted_Events_Run_2893
  • 36717 03/22/00 18:44 adev Accepted_Events_Run_2892
  • 36714 03/22/00 16:25 LeRose Beam Back/ OTR problems
  • 36712 03/22/00 16:09 Quemener Beam OFF
  • 36710 03/22/00 15:56 jpchen Shift summary
  • 36709 03/22/00 15:52 Chang Optic strip charts
  • 36708 03/22/00 15:46 adev Accepted_Events_Run_2891
  • 36706 03/22/00 15:41 Chang Beam strip charts
  • 36705 03/22/00 15:38 Chang Magnets
  • 36701 03/22/00 12:52 adev Accepted_Events_Run_2889
  • 36698 03/22/00 11:49 adev Accepted_Events_Run_2886
  • 36697 03/22/00 11:46 Adaq Raster spot size
  • 36694 03/22/00 11:37 adev Accepted_Events_Run_2885
  • 36691 03/22/00 11:32 adev Accepted_Events_Run_2884
  • 36688 03/22/00 11:28 adev Accepted_Events_Run_2883
  • 36685 03/22/00 11:16 adev Accepted_Events_Run_2882
  • 36682 03/22/00 11:06 adev Accepted_Events_Run_2881
  • 36680 03/22/00 10:47 jpchen OTR
  • 36679 03/22/00 10:19 Chang OTR
  • 36677 03/22/00 09:51 adev Accepted_Events_Run_2879
  • 36676 03/22/00 09:51 adev Accepted_Events_Run_2880
  • 36672 03/22/00 08:00 M. Rvachev Shift summary
  • 36671 03/22/00 07:50 M. Rvachev Magnets screen
  • 36670 03/22/00 07:49 M. Rvachev Magnet stripcharts
  • 36669 03/22/00 07:46 M. Rvachev Beam stripcharts
  • 36668 03/22/00 07:42 adev Accepted_Events_Run_2878
  • 36665 03/22/00 05:31 adev Accepted_Events_Run_2877
  • 36662 03/22/00 04:16 M. Rvachev List of runs for beam displacement study
  • 36660 03/22/00 03:37 adev Accepted_Events_Run_2876
  • 36657 03/22/00 03:17 adev Accepted_Events_Run_2875
  • 36655 03/22/00 03:05 M. Rvachev spot++ for runs 2873, 2875
  • 36653 03/22/00 02:36 adev Accepted_Events_Run_2874
  • 36650 03/22/00 02:29 adev Accepted_Events_Run_2873
  • 36647 03/22/00 01:10 adev Accepted_Events_Run_2870
  • 36645 03/22/00 01:00 adev Accepted_Events_Run_2869
  • 36644 03/22/00 01:00 M. Rvachev spot++ for runs 2869, 2870
  • 36640 03/22/00 00:06 adev Accepted_Events_Run_2868
  • 36637 03/22/00 00:04 adev Accepted_Events_Run_2867
  • Tuesday

  • 36633 03/21/00 23:59 Chang Shift summary
  • 36632 03/21/00 23:55 merchez beam optics stripchart
  • 36631 03/21/00 23:52 adev Accepted_Events_Run_2866
  • 36630 03/21/00 23:52 merchez magnet stripchart
  • 36628 03/21/00 23:44 Chang Harp Scan
  • 36626 03/21/00 23:31 adev Accepted_Events_Run_2865
  • 36623 03/21/00 15:58 adaq shift summary
  • 36622 03/21/00 15:49 Kelly shift summary: no beam
  • 36619 03/21/00 14:46 adev Accepted_Events_Run_2861
  • 36616 03/21/00 14:31 adev Accepted_Events_Run_2860
  • 36614 03/21/00 10:16 Adaq kin 12 magnets controls screen
  • 36612 03/21/00 08:47 G.Quemener kinematics 15 2bbu peak
  • 36611 03/21/00 08:32 Saha Denergy strip chart
  • 36610 03/21/00 08:31 Saha Magnet Strip Charts
  • 36609 03/21/00 08:29 Saha HAC screen
  • 36608 03/21/00 08:28 Saha Beam strip charts
  • 36606 03/21/00 08:01 aniol shift summary
  • 36605 03/21/00 08:01 Adaq Kin 15 Run Summary
  • 36604 03/21/00 07:54 adev Accepted_Events_Run_2855
  • 36602 03/21/00 07:30 adev Accepted_Events_Run_2854
  • 36597 03/21/00 05:48 adev Accepted_Events_Run_2853
  • 36596 03/21/00 04:13 adev Accepted_Events_Run_2852
  • 36593 03/21/00 02:36 adev Accepted_Events_Run_2851
  • 36590 03/21/00 00:57 aniol harp scan #688/#2850
  • 36587 03/21/00 00:45 Adaq harp scan
  • 36586 03/21/00 00:32 Margaziotis Shift Summary
  • 36585 03/21/00 00:29 aniol otr
  • 36582 03/21/00 00:25 adev Accepted_Events_Run_2850
  • Monday

  • 36579 03/20/00 23:05 adev Accepted_Events_Run_2849
  • 36577 03/20/00 21:17 adev Accepted_Events_Run_2848
  • 36574 03/20/00 19:34 adev Accepted_Events_Run_2847
  • 36571 03/20/00 17:55 adev Accepted_Events_Run_2846
  • 36568 03/20/00 16:48 Margaziotis Dipole strip charts
  • 36567 03/20/00 16:41 Margaziotis Magnet settings chart
  • 36566 03/20/00 16:39 Margaziotis Beam denergy chart
  • 36565 03/20/00 16:35 Margaziotis Magnet strip charts
  • 36564 03/20/00 16:33 Margaziotis Beam strip charts
  • 36563 03/20/00 16:16 adev Accepted_Events_Run_2845
  • 36559 03/20/00 15:58 Kelly shift summary
  • 36558 03/20/00 14:36 adev Accepted_Events_Run_2844
  • 36555 03/20/00 13:47 Adaq beam position strip
  • 36554 03/20/00 13:41 Kelly beam&magnet strip charts
  • 36553 03/20/00 12:55 adev Accepted_Events_Run_2843
  • 36549 03/20/00 11:10 adev Accepted_Events_Run_2842
  • 36547 03/20/00 09:58 Kelly OTR at 27 uA
  • 36545 03/20/00 09:39 adev Accepted_Events_Run_2841
  • 36542 03/20/00 07:56 aniol shift summary
  • 36540 03/20/00 06:58 adev Accepted_Events_Run_2839
  • 36537 03/20/00 05:16 adev Accepted_Events_Run_2838
  • 36534 03/20/00 03:44 Adaq magnet screen
  • 36533 03/20/00 03:42 aniol magnet strip charts
  • 36532 03/20/00 03:42 Adaq beam screen
  • 36530 03/20/00 03:26 Adaq otr
  • 36529 03/20/00 03:07 adev Accepted_Events_Run_2837
  • 36527 03/20/00 01:30 adev Accepted_Events_Run_2836
  • 36523 03/20/00 00:13 G. Quemener Shift Summary
  • 36522 03/20/00 00:03 adev Accepted_Events_Run_2835
  • Sunday

  • 36519 03/19/00 22:18 adev Accepted_Events_Run_2834
  • 36516 03/19/00 20:39 adev Accepted_Events_Run_2833
  • 36513 03/19/00 19:20 adaq intermediate kinematics 15 summary
  • 36512 03/19/00 19:06 adev Accepted_Events_Run_2832
  • 36509 03/19/00 18:42 adaq magnet strip charts
  • 36508 03/19/00 18:38 adaq beam strip charts
  • 36507 03/19/00 17:34 kemener. Bzh OTR screen
  • 36504 03/19/00 17:28 adev Accepted_Events_Run_2831
  • 36501 03/19/00 17:15 adev Accepted_Events_Run_2830
  • 36499 03/19/00 16:32 adev Accepted_Events_Run_2829
  • 36496 03/19/00 15:21 Kelly beam ooptic strip chart
  • 36495 03/19/00 15:11 Kelly magnet strip charts
  • 36493 03/19/00 14:48 bogdanw harp scan raster OFF, run 671
  • 36492 03/19/00 14:44 bogdanw harp scan raster ON, run 670
  • 36491 03/19/00 13:04 adev Accepted_Events_Run_2828
  • 36488 03/19/00 12:22 adev Accepted_Events_Run_2827
  • 36486 03/19/00 10:41 adev Accepted_Events_Run_2826
  • 36482 03/19/00 07:57 Aniol shift summary
  • 36480 03/19/00 07:27 hu magnet screen
  • 36478 03/19/00 06:57 Adaq ootr screen
  • 36477 03/19/00 00:10 Adaq Data summary this shift
  • 36476 03/19/00 00:09 Epstein Shift Summary
  • Saturday

  • 36474 03/18/00 23:12 adev Accepted_Events_Run_2824
  • 36471 03/18/00 22:46 M. Rvachev Harp scan no raster
  • 36470 03/18/00 21:55 M. Rvachev Harp scan raster on
  • 36469 03/18/00 21:43 adev Accepted_Events_Run_2823
  • 36467 03/18/00 20:52 M. Rvachev Still see a problem in VDC
  • 36464 03/18/00 18:04 M. Rvachev Magnet charts
  • 36463 03/18/00 17:59 M. Rvachev magnets screen and beam history
  • 36462 03/18/00 17:57 adev Accepted_Events_Run_2820
  • 36459 03/18/00 17:26 M. Rvachev Some info on VDC problems
  • 36458 03/18/00 17:05 Adaq Kin. 15 number of events to data
  • 36456 03/18/00 16:09 quemener shift summary
  • 36454 03/18/00 16:05 bogdanw TDC for VDC
  • 36453 03/18/00 15:39 adev Accepted_Events_Run_2817
  • 36452 03/18/00 15:39 adev Accepted_Events_Run_2817
  • 36450 03/18/00 15:35 adev Accepted_Events_Run_2818
  • 36443 03/18/00 12:46 adev Accepted_Events_Run_2816
  • 36440 03/18/00 11:52 adaq magnets strip charts
  • 36439 03/18/00 11:51 adaq beam strip charts
  • 36438 03/18/00 11:49 adaq magnets controls screen
  • 36436 03/18/00 11:26 adev Accepted_Events_Run_2815
  • 36434 03/18/00 11:13 adev Accepted_Events_Run_2814
  • 36431 03/18/00 10:49 R. Michaels re: event types
  • 36429 03/18/00 10:43 adev Accepted_Events_Run_2813
  • 36428 03/18/00 10:38 R. Michaels loss of BPM EPICs data
  • 36423 03/18/00 09:25 adev Accepted_Events_Run_2810
  • 36420 03/18/00 08:58 adev Accepted_Events_Run_2809
  • 36418 03/18/00 08:34 adev Accepted_Events_Run_2808
  • 36417 03/18/00 08:28 F.Benmokhtar e_miss and tc_cor for kin#15 (For 16.5 M)
  • 36413 03/18/00 08:06 adev Accepted_Events_Run_2807
  • 36410 03/18/00 08:04 Adaq shift summary
  • 36409 03/18/00 08:04 Adaq Kinematic 15 record.
  • 36407 03/18/00 08:03 adev Accepted_Events_Run_2806
  • 36406 03/18/00 07:26 Adaq magnet strip charts
  • 36405 03/18/00 07:22 Adaq energy strip chart
  • 36404 03/18/00 06:50 adev Accepted_Events_Run_2805
  • 36402 03/18/00 05:04 adev Accepted_Events_Run_2804
  • 36399 03/18/00 04:09 Adaq otr screen
  • 36396 03/18/00 03:58 adev Accepted_Events_Run_2803
  • 36395 03/18/00 02:50 F.Benmokhtar Energy and Beam stript charts
  • 36394 03/18/00 02:48 F.Benmokhtar Magnet stript charts.
  • 36393 03/18/00 02:38 Adaq Kinematic 15 record.
  • 36392 03/18/00 02:29 hu spot size
  • 36391 03/18/00 02:17 adev Accepted_Events_Run_2802
  • 36388 03/18/00 00:32 adev Accepted_Events_Run_2801
  • 36385 03/18/00 00:04 M. Rvachev E_miss for kin 15
  • 36384 03/18/00 00:02 Epstein Shift Summary
  • Friday

  • 36382 03/17/00 23:54 M. Rvachev 2 cards silent on hadron v1 VDC plane
  • 36381 03/17/00 22:54 adev Accepted_Events_Run_2800
  • 36377 03/17/00 20:52 Epstein OTR
  • 36376 03/17/00 20:44 M. Rvachev Harp scan with no raster
  • 36375 03/17/00 20:30 M. Rvachev Harp scan with raster on
  • 36374 03/17/00 20:23 adev Accepted_Events_Run_2799
  • 36370 03/17/00 18:39 adev Accepted_Events_Run_2798
  • 36369 03/17/00 17:33 Epstein magnet strip charts
  • 36368 03/17/00 17:31 Adaq Beam strip chart
  • 36367 03/17/00 17:27 M. Rvachev Target IOC rebooted
  • 36366 03/17/00 17:26 Epstein beam strip charts
  • 36365 03/17/00 17:25 Epstein magnets
  • 36363 03/17/00 16:55 adev Accepted_Events_Run_2797
  • 36360 03/17/00 16:00 Quemener Shift summary
  • 36359 03/17/00 15:33 Saha Arc-Integral measurement
  • 36357 03/17/00 14:59 adaq denergy problem
  • 36356 03/17/00 14:38 adev Accepted_Events_Run_2796
  • 36353 03/17/00 13:20 adev Accepted_Events_Run_2795
  • 36347 03/17/00 11:39 adev Accepted_Events_Run_2792
  • 36346 03/17/00 10:32 Margaziotis JT valve settings
  • 36345 03/17/00 10:30 Margaziotis ESR pressure
  • 36344 03/17/00 10:28 quemener magnet control screen
  • 36341 03/17/00 08:35 adev Accepted_Events_Run_2790
  • 36338 03/17/00 08:02 F.Benmokhtar Shift Summary.
  • 36337 03/17/00 07:54 adev Accepted_Events_Run_2789
  • 36335 03/17/00 07:30 adev Accepted_Events_Run_2788
  • 36332 03/17/00 06:21 Gayou cryo target JT valves setting
  • 36331 03/17/00 06:19 F.Benmokhtar Energy and Beam stript charts.
  • 36330 03/17/00 05:43 F.Benmokhtar P0_h and P0_e Stript charts
  • 36327 03/17/00 05:02 adev Accepted_Events_Run_2787
  • 36324 03/17/00 04:11 adev Accepted_Events_Run_2786
  • 36323 03/17/00 03:40 adev Accepted_Events_Run_2785
  • 36321 03/17/00 03:35 Gayou cryotarget JT valve settings
  • 36319 03/17/00 03:31 hu spot size
  • 36317 03/17/00 03:13 adev Accepted_Events_Run_2784
  • 36314 03/17/00 03:04 adev Accepted_Events_Run_2783
  • 36312 03/17/00 03:03 F.Benmokhtar Beam and Energy stript charts
  • 36310 03/17/00 02:55 adev Accepted_Events_Run_2782
  • 36306 03/17/00 02:37 adev Accepted_Events_Run_2781
  • 36305 03/17/00 02:33 F.Benmokhtar Harp Scan with raster on.
  • 36303 03/17/00 02:26 F.Benmokhtar HARP Scan Run#655
  • 36302 03/17/00 01:27 F.Benmokhtar Magnet Screen for kin#15
  • 36301 03/17/00 00:57 adev Accepted_Events_Run_2780
  • 36297 03/17/00 00:51 adev Accepted_Events_Run_2779
  • 36296 03/17/00 00:48 F.Benmokhtar Magnet screen for carbon pointing
  • 36295 03/17/00 00:40 adev Accepted_Events_Run_2778
  • 36291 03/17/00 00:05 adev Accepted_Events_Run_2777
  • Thursday

  • 36288 03/16/00 23:58 M. Rvachev Shift summary
  • 36287 03/16/00 23:21 R Roche magnet strip charts
  • 36286 03/16/00 21:51 M. Rvachev Statistics
  • 36283 03/16/00 18:00 adev Accepted_Events_Run_2775
  • 36281 03/16/00 17:27 adev Accepted_Events_Run_2774
  • 36278 03/16/00 16:09 adev Accepted_Events_Run_2773
  • 36274 03/16/00 15:54 adev Accepted_Events_Run_2771
  • 36270 03/16/00 12:03 adev Accepted_Events_Run_2770
  • 36268 03/16/00 12:00 adev Accepted_Events_Run_2769
  • 36264 03/16/00 11:32 adev Accepted_Events_Run_2766
  • Wednesday

  • 36261 03/15/00 11:58 adev Accepted_Events_Run_2765
  • 36259 03/15/00 11:57 adev Accepted_Events_Run_2764
  • 36256 03/15/00 11:54 adev Accepted_Events_Run_2763
  • 36253 03/15/00 11:34 adev Accepted_Events_Run_2762
  • 36249 03/15/00 11:24 adev Accepted_Events_Run_2761
  • 36246 03/15/00 10:07 adev Accepted_Events_Run_2760
  • 36243 03/15/00 09:21 adev Accepted_Events_Run_2759
  • 36240 03/15/00 09:20 adev Accepted_Events_Run_2758
  • Tuesday

  • 36238 03/14/00 16:25 R. Michaels tests... adaqfs
  • 36237 03/14/00 16:16 R. Michaels adaqfs
  • Monday

  • 36233 03/13/00 15:57 A. Hotta Target operation at sleeping
  • 36232 03/13/00 14:49 F.Benmokhtar Kin#12 Run Summary.
  • 36230 03/13/00 08:08 kuss target sleeps
  • 36229 03/13/00 07:43 markowitz shift summary
  • 36226 03/13/00 04:13 adev Accepted_Events_Run_2755
  • 36225 03/13/00 03:50 markowitz optics screen
  • 36224 03/13/00 03:49 markowitz magnet screen
  • 36223 03/13/00 03:48 markowitz screen shots
  • 36220 03/13/00 03:42 adev Accepted_Events_Run_2754
  • 36219 03/13/00 01:02 adev Accepted_Events_Run_2753
  • 36216 03/13/00 00:21 kuss cro target JT valve setting
  • Sunday

  • 36215 03/12/00 23:59 F.Benmokhtar Shift Summary
  • 36213 03/12/00 23:57 F.Benmokhtar Run summary for kin#12 until 00:00
  • 36212 03/12/00 22:31 F.Benmokhtar OTR screen.
  • 36209 03/12/00 22:16 adev Accepted_Events_Run_2752
  • 36208 03/12/00 22:01 Z. Chai beam energy strip plot
  • 36207 03/12/00 20:40 A. Hotta ESR cool_P_return_R high value change
  • 36206 03/12/00 19:34 adev Accepted_Events_Run_2751
  • 36203 03/12/00 18:36 Z. Chai E-Arm low helium level
  • 36202 03/12/00 18:04 z. chai energy strip plot
  • 36201 03/12/00 18:01 z. chai energy stripplot
  • 36200 03/12/00 16:59 adev Accepted_Events_Run_2750
  • 36195 03/12/00 15:58 Adaq beam optics strip chart
  • 36194 03/12/00 15:28 Epstein Shift Summary
  • 36193 03/12/00 15:08 epstein beam energy strip chart
  • 36191 03/12/00 14:26 adev Accepted_Events_Run_2749
  • 36187 03/12/00 14:22 adev Accepted_Events_Run_2748
  • 36184 03/12/00 14:14 adev Accepted_Events_Run_2747
  • 36183 03/12/00 14:07 Adaq otr screen
  • 36182 03/12/00 14:06 quemener harp scans
  • 36180 03/12/00 13:11 quemener Magnet control screen
  • 36179 03/12/00 13:11 mitchell target cooldown
  • 36178 03/12/00 09:06 Voutier & ot Kin # 12 runs record
  • 36176 03/12/00 08:15 Eric Voutier Shift Summary
  • 36175 03/12/00 07:00 mitchell target - ESR
  • 36174 03/12/00 05:05 adev Accepted_Events_Run_2746
  • 36171 03/12/00 04:04 adev Accepted_Events_Run_2745
  • 36169 03/12/00 03:38 Eric Voutier Harp scans # 639 & 640
  • 36167 03/12/00 03:18 Eric Voutier Magnet setting screen
  • 36166 03/12/00 02:55 adev Accepted_Events_Run_2744
  • 36164 03/12/00 02:24 kuss target cool-down
  • 36163 03/12/00 00:14 Epstein spectrometer fields reset
  • Saturday

  • 36161 03/11/00 23:15 kuss target cool-down
  • 36160 03/11/00 22:47 Higinbotham Cooling Back Down
  • 36159 03/11/00 20:56 kuss cryo-target warm-up
  • 36158 03/11/00 20:44 bogdanw magnets
  • 36157 03/11/00 20:42 bogdanw current
  • 36156 03/11/00 20:38 Epstein spectrometer shut down
  • 36155 03/11/00 20:21 kuss cryo-target warm-up
  • 36154 03/11/00 19:53 Epstein ESR CAMAC card swap
  • 36153 03/11/00 19:50 adev Accepted_Events_Run_2743
  • 36149 03/11/00 17:53 bogdanw OTR
  • 36148 03/11/00 17:52 Z. Chai harp scan 635
  • 36147 03/11/00 17:44 Z. Chai harp scan 634
  • 36146 03/11/00 17:34 F.Benmokhtar e_miss%reactez for kin#3
  • 36145 03/11/00 17:31 adev Accepted_Events_Run_2742
  • 36143 03/11/00 17:06 F.Benmokhtar e_miss for kin#12
  • 36142 03/11/00 16:30 F.Benmokhtar Shif Summary
  • 36141 03/11/00 16:28 F.Benmokhtar Run S ummary for Kin#12 until 16:00
  • 36139 03/11/00 15:51 F.Benmokhtar Energy stript chart for the end of the shift
  • 36138 03/11/00 15:49 F.Benmokhtar Beam stript chart for the end of the shift
  • 36137 03/11/00 15:45 adev Accepted_Events_Run_2741
  • 36133 03/11/00 13:11 adev Accepted_Events_Run_2740
  • 36131 03/11/00 12:52 adev Accepted_Events_Run_2739
  • 36128 03/11/00 12:27 chang otr check for shiftn checklist
  • 36127 03/11/00 10:29 F.Benmokhtar Accoustical alarm in the Right gasmaster gauge
  • 36126 03/11/00 10:24 G. Chang Energy Stript Chart
  • 36125 03/11/00 10:22 adev Accepted_Events_Run_2738
  • 36122 03/11/00 10:15 F.Benmokhtar E89044_beam stript chart
  • 36121 03/11/00 08:08 Voutier & ot Kin # 12 record
  • 36118 03/11/00 06:52 adev Accepted_Events_Run_2737
  • 36116 03/11/00 06:06 adev Accepted_Events_Run_2736
  • 36112 03/11/00 04:03 markowitz energy width
  • 36111 03/11/00 03:47 Adaq magnets controls screen
  • 36110 03/11/00 03:28 Adaq otr
  • 36108 03/11/00 03:21 adev Accepted_Events_Run_2735
  • 36106 03/11/00 03:09 adev Accepted_Events_Run_2734
  • Friday

  • 36103 03/10/00 23:53 Eric Voutier Shift Summary
  • 36100 03/10/00 16:03 kriswhniw Shift summary
  • 36098 03/10/00 15:37 M Jones change to epicsLogger
  • 36097 03/10/00 15:28 R. Michaels beam ?
  • 36096 03/10/00 08:38 krishniw Energy lock
  • 36095 03/10/00 08:33 R. Michaels beam down time
  • 36094 03/10/00 08:21 F.Benmokhtar tc_cor for kin#12
  • 36093 03/10/00 08:16 F.Benmokhtar 2bbu rate for Kin#12
  • 36091 03/10/00 08:00 markowitz summary for kinematics 12
  • 36090 03/10/00 07:59 markowitz shift summary
  • 36089 03/10/00 07:41 adev Accepted_Events_Run_2733
  • 36086 03/10/00 06:53 adev Accepted_Events_Run_2732
  • 36083 03/10/00 06:14 markowitz strip charts
  • 36081 03/10/00 05:39 adev Accepted_Events_Run_2729
  • 36079 03/10/00 04:00 Adaq beam
  • 36077 03/10/00 03:39 markowitz magnets
  • 36076 03/10/00 03:34 F.Benmokhtar Harp Scan For kin#12 (Raster off)
  • 36074 03/10/00 03:33 adev Accepted_Events_Run_2728
  • 36072 03/10/00 03:26 F.Benmokhtar Harp Scan for kin#12 (Raster on)
  • 36071 03/10/00 03:16 adev Accepted_Events_Run_2727
  • 36068 03/10/00 03:04 adev Accepted_Events_Run_2726
  • 36065 03/10/00 01:08 adev Accepted_Events_Run_2725
  • 36062 03/10/00 01:07 adev Accepted_Events_Run_2724
  • 36059 03/10/00 00:26 F.Benmokhtar Prescalers changed
  • 36057 03/10/00 00:04 Quemener et Summary for kinemtics #12
  • 36055 03/10/00 00:01 G. Quemener Shift Summary
  • Thursday

  • 36053 03/09/00 23:55 adev Accepted_Events_Run_2722
  • 36050 03/09/00 23:18 adev Accepted_Events_Run_2721
  • 36048 03/09/00 23:00 adev Accepted_Events_Run_2720
  • 36044 03/09/00 22:42 adev Accepted_Events_Run_2719
  • 36041 03/09/00 22:21 adev Accepted_Events_Run_2718
  • 36040 03/09/00 22:18 mitchell ep energy measurement
  • 36037 03/09/00 22:04 adev Accepted_Events_Run_2717
  • 36034 03/09/00 21:44 adev Accepted_Events_Run_2716
  • 36031 03/09/00 21:23 adev Accepted_Events_Run_2715
  • 36030 03/09/00 21:20 quemener spectrometer motion
  • 36029 03/09/00 21:11 Adaq magnets controls screen
  • 36028 03/09/00 20:42 Perdrisat & Kin # 9 running log
  • 36025 03/09/00 20:26 adev Accepted_Events_Run_2714
  • 36022 03/09/00 20:17 adev Accepted_Events_Run_2713
  • 36021 03/09/00 19:18 adev Accepted_Events_Run_2712
  • 36018 03/09/00 18:16 adev Accepted_Events_Run_2711
  • 36015 03/09/00 18:03 quemener beam strip tool
  • 36014 03/09/00 17:48 quemener magnets strip charts
  • 36013 03/09/00 17:48 quemener magnets strip charts
  • 36012 03/09/00 17:15 adev Accepted_Events_Run_2710
  • 36009 03/09/00 16:14 adev Accepted_Events_Run_2709
  • 36005 03/09/00 15:54 Eric Voutier Shift summary
  • 36004 03/09/00 15:28 Eric Voutier Magnet strip chart
  • 36003 03/09/00 15:26 Eric Voutier Beam strip chart
  • 36000 03/09/00 15:00 adev Accepted_Events_Run_2708
  • 35998 03/09/00 14:17 adev Accepted_Events_Run_2707
  • 35997 03/09/00 14:10 R. Michaels roc14 crash
  • 35994 03/09/00 13:50 folts Q1 failure
  • 35992 03/09/00 13:20 Eric Voutier Harp scan #626 & #627
  • 35991 03/09/00 12:54 adev Accepted_Events_Run_2705
  • 35989 03/09/00 12:53 R. Michaels beam spot
  • 35987 03/09/00 12:51 adev Accepted_Events_Run_2704
  • 35984 03/09/00 12:48 R. Michaels beam back, harp scans
  • 35982 03/09/00 12:48 adev Accepted_Events_Run_2703
  • 35980 03/09/00 10:33 Lindgren Access, Raster
  • 35979 03/09/00 09:58 Lindgren, Vo Cathode, beam, cesiate
  • 35978 03/09/00 08:10 adev Accepted_Events_Run_2702
  • 35975 03/09/00 07:57 M. Rvachev Shift summary
  • 35974 03/09/00 07:38 adev Accepted_Events_Run_2701
  • 35971 03/09/00 07:04 M. Rvachev List of all runs taken so far(2)
  • 35970 03/09/00 06:47 M. Rvachev List of runs taken so far
  • 35968 03/09/00 05:53 adev Accepted_Events_Run_2700
  • 35966 03/09/00 05:36 F.Benmokhtar e_miss and tc_cor for kin#9.
  • 35965 03/09/00 04:37 M. Rvachev Beam history and magnets screen
  • 35964 03/09/00 04:35 M. Rvachev Magnet stripcharts
  • 35963 03/09/00 04:35 adev Accepted_Events_Run_2699
  • 35960 03/09/00 03:07 adev Accepted_Events_Run_2698
  • 35957 03/09/00 01:43 adev Accepted_Events_Run_2697
  • 35954 03/09/00 00:18 adev Accepted_Events_Run_2696
  • 35950 03/09/00 00:00 LeRose Shift summary
  • Wednesday

  • 35949 03/08/00 23:51 kuss gaussian integration (was: summary kin #9)
  • 35948 03/08/00 23:39 kuss new espace and the width of kin#3
  • 35947 03/08/00 23:25 LeRose PMT HV GUI all zeroes
  • 35946 03/08/00 23:02 kuss summary kin #9
  • 35945 03/08/00 22:53 adev Accepted_Events_Run_2695
  • 35942 03/08/00 22:19 kuss summary kin #9
  • 35941 03/08/00 21:33 adev Accepted_Events_Run_2694
  • 35938 03/08/00 20:45 LeRose Beam Strip Charts III
  • 35937 03/08/00 20:44 LeRose Beam Strip Charts II
  • 35936 03/08/00 20:40 LeRose Beam stripchart
  • 35935 03/08/00 20:34 bogdanw magnet stripcharts
  • 35934 03/08/00 20:11 adev Accepted_Events_Run_2693
  • 35931 03/08/00 19:01 LeRose Beam Strip chart
  • 35930 03/08/00 18:59 LeRose Magnet Strip Charts
  • 35929 03/08/00 18:52 adev Accepted_Events_Run_2692
  • 35926 03/08/00 18:03 kuss estimate for kin#9
  • 35925 03/08/00 18:00 LeRose Beam Back
  • 35923 03/08/00 17:17 LeRose Beam OFF
  • 35922 03/08/00 17:09 adev Accepted_Events_Run_2691
  • 35920 03/08/00 16:55 LeRose MCC communication
  • 35919 03/08/00 16:32 LeRose HV was off
  • 35916 03/08/00 16:24 adev Accepted_Events_Run_2690
  • 35915 03/08/00 16:18 LeRose T1 T2 not counting
  • 35914 03/08/00 16:08 Krishniw Shift summary
  • 35911 03/08/00 15:57 adev Accepted_Events_Run_2689
  • 35908 03/08/00 15:11 adev Accepted_Events_Run_2688
  • 35906 03/08/00 14:36 A.Hotta He-3 Target
  • 35903 03/08/00 13:45 adev Accepted_Events_Run_2687
  • 35902 03/08/00 12:28 adev Accepted_Events_Run_2686
  • 35899 03/08/00 11:09 adev Accepted_Events_Run_2685
  • 35896 03/08/00 10:51 krishniw Magnetic screen
  • 35895 03/08/00 10:50 krishniw beam strip chart
  • 35894 03/08/00 09:46 adev Accepted_Events_Run_2684
  • 35891 03/08/00 08:53 Nilanga Emiss resolution at high density
  • 35889 03/08/00 08:00 Eric Voutier Shift summary
  • 35887 03/08/00 07:23 adev Accepted_Events_Run_2682
  • 35884 03/08/00 06:05 adev Accepted_Events_Run_2681
  • 35881 03/08/00 04:47 adev Accepted_Events_Run_2680
  • 35880 03/08/00 04:46 Eric Voutier Harp scan # 623
  • 35876 03/08/00 02:37 adev Accepted_Events_Run_2679
  • 35873 03/08/00 02:20 adev Accepted_Events_Run_2678
  • 35871 03/08/00 01:51 finn baeam strip charts
  • 35870 03/08/00 01:50 finn beam strip charts
  • 35869 03/08/00 01:36 finn magnet strip chart
  • 35868 03/08/00 01:32 finn magnet summary display
  • 35867 03/08/00 01:00 adev Accepted_Events_Run_2677
  • Tuesday

  • 35863 03/07/00 23:55 Perdrisat 16-24 shift record
  • 35861 03/07/00 23:39 adev Accepted_Events_Run_2676
  • 35858 03/07/00 23:29 adev Accepted_Events_Run_2675
  • 35855 03/07/00 23:18 Perdrisat H-gas alarm
  • 35854 03/07/00 23:13 adev Accepted_Events_Run_2674
  • 35852 03/07/00 22:43 Punjabi E-miss spectra
  • 35850 03/07/00 22:36 adev Accepted_Events_Run_2673
  • 35847 03/07/00 22:27 adev Accepted_Events_Run_2672
  • 35844 03/07/00 21:39 Saha Summary of Kinematics #3
  • 35843 03/07/00 21:37 adev Accepted_Events_Run_2671
  • 35840 03/07/00 21:30 adev Accepted_Events_Run_2670
  • 35837 03/07/00 21:17 adev Accepted_Events_Run_2669
  • 35834 03/07/00 20:44 adev Accepted_Events_Run_2668
  • 35831 03/07/00 20:09 adev Accepted_Events_Run_2667
  • 35827 03/07/00 19:33 adev Accepted_Events_Run_2666
  • 35826 03/07/00 18:59 adev Accepted_Events_Run_2665
  • 35822 03/07/00 18:37 adev Accepted_Events_Run_2664
  • 35820 03/07/00 18:29 Perdrisat hspect and espect
  • 35819 03/07/00 18:27 Perdrisat magnets and e89044_beam
  • 35817 03/07/00 18:10 Perdrisat harp scan, FFB on
  • 35815 03/07/00 18:08 adev Accepted_Events_Run_2663
  • 35813 03/07/00 17:31 Nilanga Emiss resolution with low density
  • 35812 03/07/00 17:17 adev Accepted_Events_Run_2662
  • 35810 03/07/00 17:06 Nilanga Emiss resolution for 12C
  • 35809 03/07/00 17:05 Nilanga Emiss resolution for 12C
  • 35807 03/07/00 16:53 adev Accepted_Events_Run_2661
  • 35804 03/07/00 16:51 kuss target density back to maximum
  • 35803 03/07/00 16:51 adev Accepted_Events_Run_2660
  • 35801 03/07/00 16:02 Saha Shift Summary, 08:00-16:00, March 7, 2000
  • 35799 03/07/00 15:58 Saha Beamline Stripchart
  • 35798 03/07/00 15:57 Saha Harp scan - No FFB, No raster
  • 35796 03/07/00 15:53 adev Accepted_Events_Run_2659
  • 35793 03/07/00 15:41 adev Accepted_Events_Run_2658
  • 35791 03/07/00 15:37 Saha Magnet stripcharts
  • 35790 03/07/00 15:36 Saha Energy locks in halls A and C
  • 35788 03/07/00 15:11 adev Accepted_Events_Run_2657
  • 35786 03/07/00 14:57 Benmokhtar Half Density
  • 35784 03/07/00 14:06 F.Benmokhtar e_miss spectra for kin3 and 6 (half and full density)
  • 35782 03/07/00 13:37 adev Accepted_Events_Run_2656
  • 35781 03/07/00 13:33 Saha HAC Medm Screen - Kin#3
  • 35779 03/07/00 12:55 Saha Hall A arc integral-Old Integral value
  • 35778 03/07/00 12:44 Saha Hall A Arc - Bdl Measurement
  • 35776 03/07/00 12:31 adev Accepted_Events_Run_2655
  • 35774 03/07/00 09:51 A. Hotta Energy graph
  • 35773 03/07/00 09:34 kuss target on empty
  • 35772 03/07/00 09:29 kuss target density change
  • 35770 03/07/00 08:04 Quemener Shift Summary
  • 35769 03/07/00 07:59 adev Accepted_Events_Run_2654
  • 35767 03/07/00 07:13 adev Accepted_Events_Run_2653
  • 35762 03/07/00 06:09 adev Accepted_Events_Run_2652
  • 35760 03/07/00 05:12 adev Accepted_Events_Run_2651
  • 35758 03/07/00 04:21 adev Accepted_Events_Run_2650
  • 35755 03/07/00 03:22 adev Accepted_Events_Run_2649
  • 35752 03/07/00 02:25 adev Accepted_Events_Run_2648
  • 35749 03/07/00 01:54 Adaq CODA rate
  • 35747 03/07/00 00:58 adev Accepted_Events_Run_2647
  • 35745 03/07/00 00:41 Finn Optics Strip charts
  • 35744 03/07/00 00:39 Finn magnet strip charts
  • 35743 03/07/00 00:28 adev Accepted_Events_Run_2646
  • Monday

  • 35738 03/06/00 23:58 Eric Voutier Shift summary
  • 35736 03/06/00 23:52 adev Accepted_Events_Run_2645
  • 35733 03/06/00 23:30 adev Accepted_Events_Run_2644
  • 35730 03/06/00 23:13 adev Accepted_Events_Run_2643
  • 35727 03/06/00 23:01 adev Accepted_Events_Run_2642
  • 35725 03/06/00 22:02 adev Accepted_Events_Run_2641
  • 35722 03/06/00 21:53 Doug Bdl change
  • 35721 03/06/00 21:44 Eric Voutier Kin#6 rate estimate
  • 35720 03/06/00 21:20 adaq OTR Check
  • 35718 03/06/00 21:07 adev Accepted_Events_Run_2640
  • 35716 03/06/00 20:44 R Roche Beam optics strip chart
  • 35715 03/06/00 20:42 R Roche Magnet Strip Charts
  • 35713 03/06/00 20:05 adev Accepted_Events_Run_2639
  • 35711 03/06/00 19:21 Eric Voutier Scan# 617 and 618
  • 35708 03/06/00 18:41 adev Accepted_Events_Run_2638
  • 35705 03/06/00 18:38 adev Accepted_Events_Run_2637
  • 35703 03/06/00 18:35 adev Accepted_Events_Run_2636
  • 35700 03/06/00 17:30 adev Accepted_Events_Run_2635
  • 35697 03/06/00 16:02 Templon shift summary
  • 35695 03/06/00 15:58 Templon harp scan / raster ON
  • 35694 03/06/00 15:58 Templon OTR screen
  • 35693 03/06/00 15:48 Templon harp scans / raster off
  • 35692 03/06/00 14:12 bogdanw beam on BeO
  • 35690 03/06/00 12:37 LeRose Back to 417 MeV/c
  • 35689 03/06/00 12:30 LeRose Electron 374 MeV/c Proc
  • 35688 03/06/00 12:02 adev Accepted_Events_Run_2634
  • 35687 03/06/00 11:30 LeRose Electron Dipole Setting
  • 35684 03/06/00 08:10 M. Rvachev Beam history and magnets screen
  • 35682 03/06/00 07:59 M. Rvachev Shift summary
  • 35681 03/06/00 07:56 M. Rvachev New trigger downloaded
  • 35680 03/06/00 04:01 M. Rvachev ED reset
  • 35679 03/06/00 03:03 M. Rvachev Problems setting momentum in ED to 417MeV
  • 35678 03/06/00 00:56 M. Rvachev Run summary for kin 94
  • 35677 03/06/00 00:51 finn more beam charts
  • 35676 03/06/00 00:50 finn beam current and position charts
  • 35675 03/06/00 00:33 adev Accepted_Events_Run_2631
  • 35673 03/06/00 00:18 M. Rvachev Beam history and magnets screen
  • 35672 03/06/00 00:01 Quemener shift and data summary
  • Sunday

  • 35669 03/05/00 23:44 adev Accepted_Events_Run_2630
  • 35666 03/05/00 20:02 adev Accepted_Events_Run_2629
  • 35663 03/05/00 19:42 adev Accepted_Events_Run_2628
  • 35661 03/05/00 19:17 quemener coda reboot
  • 35659 03/05/00 18:20 adev Accepted_Events_Run_2626
  • 35656 03/05/00 17:35 adev Accepted_Events_Run_2625
  • 35652 03/05/00 16:46 adev Accepted_Events_Run_2624
  • 35649 03/05/00 16:14 adev Accepted_Events_Run_2623
  • 35647 03/05/00 16:12 Eric Voutier Shift Summary
  • 35646 03/05/00 16:09 Eric Voutier Harp Scan
  • 35643 03/05/00 15:57 Eric Voutier Kin#94 settings
  • 35642 03/05/00 15:44 adev Accepted_Events_Run_2622
  • 35640 03/05/00 15:22 adev Accepted_Events_Run_2621
  • 35637 03/05/00 14:18 Saha Arc Energy Measurment, 2 pass, 1.254 GeV
  • 35636 03/05/00 13:24 Saha Harp Scan , no raster, Size unacceptable
  • 35635 03/05/00 13:13 Saha OTR screen, 5 uA, Achromatic mode
  • 35632 03/05/00 12:39 adev Accepted_Events_Run_2620
  • 35630 03/05/00 11:54 adev Accepted_Events_Run_2619
  • 35627 03/05/00 11:31 adev Accepted_Events_Run_2618
  • 35623 03/05/00 11:04 adev Accepted_Events_Run_2617
  • 35621 03/05/00 10:00 Eric Voutier Kin#93 rate estimate
  • 35620 03/05/00 09:25 adev Accepted_Events_Run_2616
  • 35617 03/05/00 09:16 adev Accepted_Events_Run_2615
  • 35615 03/05/00 08:59 adev Accepted_Events_Run_2614
  • 35612 03/05/00 08:37 adev Accepted_Events_Run_2613
  • 35608 03/05/00 07:57 M. Rvachev Espace imperfections
  • 35607 03/05/00 07:57 Finn run summary 03/05/00 00:00 to 08:00
  • 35605 03/05/00 07:38 finn magnet summary screen
  • 35604 03/05/00 07:24 adev Accepted_Events_Run_2612
  • 35602 03/05/00 06:58 adev Accepted_Events_Run_2611
  • 35599 03/05/00 06:48 adev Accepted_Events_Run_2610
  • 35596 03/05/00 06:36 adev Accepted_Events_Run_2609
  • 35593 03/05/00 06:13 adev Accepted_Events_Run_2608
  • 35590 03/05/00 05:50 adev Accepted_Events_Run_2607
  • 35587 03/05/00 05:34 adev Accepted_Events_Run_2606
  • 35583 03/05/00 05:08 adev Accepted_Events_Run_2605
  • 35580 03/05/00 04:41 adev Accepted_Events_Run_2604
  • 35578 03/05/00 04:32 finn magnet summary screen
  • 35576 03/05/00 04:14 adev Accepted_Events_Run_2603
  • 35573 03/05/00 03:52 M. Rvachev Please don't run netscape on adaqs2
  • 35572 03/05/00 03:51 adev Accepted_Events_Run_2602
  • 35569 03/05/00 03:38 adev Accepted_Events_Run_2601
  • 35566 03/05/00 03:33 adev Accepted_Events_Run_2600
  • 35563 03/05/00 03:29 adev Accepted_Events_Run_2599
  • 35560 03/05/00 03:24 adev Accepted_Events_Run_2598
  • 35557 03/05/00 03:20 adev Accepted_Events_Run_2597
  • 35554 03/05/00 03:14 finn magnet screen
  • 35553 03/05/00 02:36 finn magnet strip chartsip
  • 35552 03/05/00 02:32 adev Accepted_Events_Run_2596
  • 35549 03/05/00 02:30 adev Accepted_Events_Run_2595
  • 35546 03/05/00 02:27 adev Accepted_Events_Run_2594
  • 35543 03/05/00 02:25 adev Accepted_Events_Run_2593
  • 35542 03/05/00 02:24 adev Accepted_Events_Run_2592
  • 35537 03/05/00 01:52 adev Accepted_Events_Run_2591
  • 35534 03/05/00 01:33 adev Accepted_Events_Run_2590
  • 35532 03/05/00 01:19 finn Datacube OTR screen frozen
  • 35527 03/05/00 01:00 adev Accepted_Events_Run_2588
  • 35523 03/05/00 00:46 adev Accepted_Events_Run_2587
  • 35522 03/05/00 00:44 finn magnet summary screen
  • 35519 03/05/00 00:33 adev Accepted_Events_Run_2586
  • 35517 03/05/00 00:30 adev Accepted_Events_Run_2585
  • 35515 03/05/00 00:28 adev Accepted_Events_Run_2584
  • 35511 03/05/00 00:24 adev Accepted_Events_Run_2583
  • 35508 03/05/00 00:15 adev Accepted_Events_Run_2582
  • 35506 03/05/00 00:13 F.Benmokhtar Shift Summary
  • 35504 03/05/00 00:02 adev Accepted_Events_Run_2581
  • Saturday

  • 35502 03/04/00 23:51 adev Accepted_Events_Run_2580
  • 35498 03/04/00 23:44 adev Accepted_Events_Run_2579
  • 35496 03/04/00 23:35 F.Benmokhtar HARP scan with raster on
  • 35495 03/04/00 23:25 adev Accepted_Events_Run_2578
  • 35491 03/04/00 23:07 F.Benmokhtar HARP scan analysis
  • 35490 03/04/00 22:55 F.Benmokhtar Harp scan screen
  • 35489 03/04/00 22:23 F.Benmokhtar Magnet Setting for kin90
  • 35488 03/04/00 21:47 F.Benmokhtar OTR Screen
  • 35487 03/04/00 20:21 adev Accepted_Events_Run_2577
  • 35484 03/04/00 20:15 adev Accepted_Events_Run_2576
  • 35481 03/04/00 20:02 adev Accepted_Events_Run_2575
  • 35479 03/04/00 19:51 adev Accepted_Events_Run_2574
  • 35475 03/04/00 19:26 F.Benmokhtar Magnet Setting for kin 81
  • 35474 03/04/00 19:17 adev Accepted_Events_Run_2573
  • 35472 03/04/00 18:48 Meekins Magnet settings for Setting 80
  • 35470 03/04/00 18:39 adev Accepted_Events_Run_2572
  • 35466 03/04/00 18:16 adev Accepted_Events_Run_2571
  • 35464 03/04/00 18:11 Meekins Magnet screens for setting 78
  • 35463 03/04/00 18:01 adev Accepted_Events_Run_2570
  • 35460 03/04/00 17:46 adev Accepted_Events_Run_2569
  • 35456 03/04/00 17:24 adev Accepted_Events_Run_2568
  • 35454 03/04/00 17:19 F.Benmokhtar Magnet setting of kin76
  • 35453 03/04/00 17:06 adev Accepted_Events_Run_2567
  • 35451 03/04/00 16:55 F.Benmokhtar Magnet setting for kin74/75
  • 35449 03/04/00 16:37 adev Accepted_Events_Run_2566
  • 35446 03/04/00 16:25 adev Accepted_Events_Run_2565
  • 35442 03/04/00 16:08 adev Accepted_Events_Run_2564
  • 35440 03/04/00 16:03 quemener Shift summary
  • 35438 03/04/00 16:01 adev Accepted_Events_Run_2563
  • 35435 03/04/00 15:47 adev Accepted_Events_Run_2562
  • 35434 03/04/00 15:44 Adaq magnets controls screen
  • 35431 03/04/00 15:24 adev Accepted_Events_Run_2561
  • 35428 03/04/00 14:57 adev Accepted_Events_Run_2560
  • 35424 03/04/00 14:15 adev Accepted_Events_Run_2559
  • 35422 03/04/00 14:02 adev Accepted_Events_Run_2558
  • 35419 03/04/00 13:32 quemener magnets controls screen
  • 35418 03/04/00 13:20 Adaq magnets controls screen
  • 35417 03/04/00 13:10 adev Accepted_Events_Run_2557
  • 35416 03/04/00 13:10 Adaq magnets controls screen
  • 35413 03/04/00 12:46 adev Accepted_Events_Run_2556
  • 35410 03/04/00 12:13 adev Accepted_Events_Run_2555
  • 35407 03/04/00 11:59 quemener magnet control screen
  • 35406 03/04/00 11:48 adev Accepted_Events_Run_2554
  • 35403 03/04/00 11:35 quemener magnets controls screen
  • 35402 03/04/00 11:23 adev Accepted_Events_Run_2553
  • 35399 03/04/00 11:03 quemener magnets controls window
  • 35398 03/04/00 10:37 adev Accepted_Events_Run_2552
  • 35395 03/04/00 10:12 quemener magnets control window
  • 35394 03/04/00 10:11 quemener magnets strip chart
  • 35393 03/04/00 09:58 adev Accepted_Events_Run_2551
  • 35389 03/04/00 09:36 adev Accepted_Events_Run_2550
  • 35387 03/04/00 09:20 adev Accepted_Events_Run_2549
  • 35384 03/04/00 09:04 adev Accepted_Events_Run_2548
  • 35380 03/04/00 08:44 adev Accepted_Events_Run_2547
  • 35378 03/04/00 08:20 adev Accepted_Events_Run_2546
  • 35375 03/04/00 08:01 Eric Voutier Eric Voutier
  • 35373 03/04/00 07:55 adev Accepted_Events_Run_2545
  • 35370 03/04/00 07:45 adev Accepted_Events_Run_2544
  • 35367 03/04/00 07:19 adev Accepted_Events_Run_2543
  • 35365 03/04/00 07:08 adev Accepted_Events_Run_2542
  • 35361 03/04/00 06:57 adev Accepted_Events_Run_2541
  • 35359 03/04/00 06:53 Eric Voutier New magnet settings
  • 35357 03/04/00 06:40 adev Accepted_Events_Run_2540
  • 35354 03/04/00 06:22 adev Accepted_Events_Run_2539
  • 35351 03/04/00 05:41 adev Accepted_Events_Run_2538
  • 35349 03/04/00 05:08 adev Accepted_Events_Run_2537
  • 35346 03/04/00 04:54 M. Rvachev Magnets screen
  • 35345 03/04/00 04:51 Eric Voutier Raster size check
  • 35344 03/04/00 04:46 Eric Voutier Beam spot aspect ratio
  • 35342 03/04/00 04:40 finn magnet stripchart
  • 35341 03/04/00 04:36 finn beam stripchart
  • 35340 03/04/00 04:32 Voutier OTR check
  • 35339 03/04/00 02:47 Saha Arc Energy Measurment, March 4, 2000, 02:30
  • 35337 03/04/00 00:12 Saha OTR at 5 uA
  • Friday

  • 35336 03/03/00 23:53 x. jiang shift activity, swing March 3rd.
  • 35335 03/03/00 23:52 Saha OTR at 5 uA
  • 35334 03/03/00 23:44 x. jiang beam current in the last few hours
  • 35333 03/03/00 23:37 Saha OTR screen
  • 35332 03/03/00 23:30 A.Hotta Kin. #66 C-12
  • 35331 03/03/00 23:29 adev Accepted_Events_Run_2536
  • 35330 03/03/00 23:28 Z. Chai e.dp_kin comparation
  • 35328 03/03/00 23:22 Z. Chai e.dp_kin comparation
  • 35325 03/03/00 23:10 adev Accepted_Events_Run_2534
  • 35322 03/03/00 23:01 adev Accepted_Events_Run_2533
  • 35319 03/03/00 22:58 adev Accepted_Events_Run_2532
  • 35316 03/03/00 22:41 Z. Chai e.dp_kin
  • 35315 03/03/00 22:40 adev Accepted_Events_Run_2531
  • 35313 03/03/00 22:34 x. jiang run 2531, high dead time at the end.
  • 35312 03/03/00 22:34 adev Accepted_Events_Run_2530
  • 35309 03/03/00 22:27 x. jiang magnet screen at run 2530. c12 elastic
  • 35306 03/03/00 22:18 Z. Chai Harp Scan 568
  • 35305 03/03/00 22:09 adev Accepted_Events_Run_2529
  • 35304 03/03/00 22:07 Z. Chai Harp Scan 567
  • 35302 03/03/00 21:57 Z. Chai Harp Scan 566
  • 35301 03/03/00 21:53 Z. Chai Harp Scan 565
  • 35300 03/03/00 21:43 Z. Chai Harp Scan 564
  • 35299 03/03/00 21:30 Z. Chai Harp Scan 563
  • 35298 03/03/00 21:13 mitchell ep energy measurement
  • 35297 03/03/00 20:16 Z. Chai Harp Scan 562
  • 35296 03/03/00 20:12 Z. Chai Harp Scan 560
  • 35295 03/03/00 20:00 x. jiang otr reading
  • 35294 03/03/00 19:51 adev Accepted_Events_Run_2528
  • 35292 03/03/00 19:36 x. jiang magnet screen before run 2528
  • 35290 03/03/00 19:27 adev Accepted_Events_Run_2527
  • 35288 03/03/00 19:22 Z. Chai raster size
  • 35286 03/03/00 19:17 adev Accepted_Events_Run_2526
  • 35283 03/03/00 19:06 adev Accepted_Events_Run_2525
  • 35280 03/03/00 19:04 adev Accepted_Events_Run_2524
  • 35277 03/03/00 18:32 adev Accepted_Events_Run_2523
  • 35275 03/03/00 18:30 Z. Chai spot check
  • 35273 03/03/00 17:57 adev Accepted_Events_Run_2522
  • 35271 03/03/00 17:56 adev Accepted_Events_Run_2521
  • 35267 03/03/00 17:49 adev Accepted_Events_Run_2520
  • 35265 03/03/00 17:06 adev Accepted_Events_Run_2519
  • 35262 03/03/00 16:31 F.Benmokhtar 8:00-16:00 Summary
  • 35259 03/03/00 08:02 Quemener No beam shift summary
  • 35255 03/03/00 06:01 adev Accepted_Events_Run_2518
  • Thursday

  • 35253 03/02/00 23:27 Eric Voutier Shift summary
  • 35250 03/02/00 15:39 kuss TWEAKING OF HRSE Q1 CURRENT
  • 35249 03/02/00 15:03 kuss setting for kin #60out
  • 35248 03/02/00 12:32 tamae strip tool
  • 35247 03/02/00 12:28 Saha Shift Summary, 08:00-12:00, March 2, 2000
  • 35246 03/02/00 12:27 Saha Elastic Kinematics for 644 MeV
  • 35245 03/02/00 12:24 tamae beam parameter
  • 35244 03/02/00 12:17 kuss summary of kin #29 at pmiss=1GeV/c
  • 35242 03/02/00 12:08 Saha Summary: Kin #29
  • 35241 03/02/00 12:02 adev Accepted_Events_Run_2517
  • 35239 03/02/00 11:55 adev Accepted_Events_Run_2516
  • 35236 03/02/00 11:52 kuss plot for #2516
  • 35235 03/02/00 11:51 kuss spot for 2516
  • 35233 03/02/00 11:50 kuss spot for #2515
  • 35232 03/02/00 11:46 adev Accepted_Events_Run_2515
  • 35230 03/02/00 11:41 Saha HAC screen, Dummy 10cm
  • 35228 03/02/00 11:33 Adaq OTR picture
  • 35225 03/02/00 11:12 adev Accepted_Events_Run_2514
  • 35223 03/02/00 10:27 adev Accepted_Events_Run_2513
  • 35222 03/02/00 09:19 Saha BdL measurement
  • 35219 03/02/00 08:48 adev Accepted_Events_Run_2512
  • 35218 03/02/00 08:32 Saha Main Hac screen, Kin #29
  • 35217 03/02/00 08:31 Saha magnet striptool
  • 35216 03/02/00 08:29 saha beam striptool
  • 35214 03/02/00 08:02 M. Rvachev Shift summary
  • 35213 03/02/00 07:42 C. Furget Number of 2bd for kin 29
  • 35212 03/02/00 07:30 adev Accepted_Events_Run_2511
  • 35209 03/02/00 06:23 adev Accepted_Events_Run_2510
  • 35205 03/02/00 04:57 adev Accepted_Events_Run_2509
  • 35202 03/02/00 04:02 adev Accepted_Events_Run_2506
  • 35199 03/02/00 02:44 adev Accepted_Events_Run_2505
  • 35196 03/02/00 01:31 adev Accepted_Events_Run_2504
  • 35193 03/02/00 00:55 C.Furget emiss plot for kin 29
  • 35192 03/02/00 00:52 C.Furget Emiss for kin 29
  • 35191 03/02/00 00:12 adev Accepted_Events_Run_2503
  • 35188 03/02/00 00:03 quemener shift summary
  • Wednesday

  • 35185 03/01/00 23:04 adev Accepted_Events_Run_2502
  • 35182 03/01/00 21:57 adev Accepted_Events_Run_2501
  • 35179 03/01/00 21:11 adev Accepted_Events_Run_2500
  • 35176 03/01/00 20:31 LeRose Magnet strip charts
  • 35175 03/01/00 20:26 adev Accepted_Events_Run_2499
  • 35172 03/01/00 20:21 LeRose Beam Strip Tool
  • 35171 03/01/00 19:59 quemener Harp scans
  • 35169 03/01/00 19:32 adev Accepted_Events_Run_2498
  • 35165 03/01/00 19:10 adev Accepted_Events_Run_2497
  • 35162 03/01/00 19:00 adev Accepted_Events_Run_2496
  • 35159 03/01/00 18:49 adev Accepted_Events_Run_2495
  • 35157 03/01/00 18:35 quemener harp scans
  • 35155 03/01/00 18:34 adev Accepted_Events_Run_2494
  • 35152 03/01/00 18:24 adev Accepted_Events_Run_2493
  • 35150 03/01/00 18:13 adaq prescales
  • 35148 03/01/00 17:04 adev Accepted_Events_Run_2492
  • 35146 03/01/00 16:48 adev Accepted_Events_Run_2491
  • 35143 03/01/00 16:38 adev Accepted_Events_Run_2490
  • 35140 03/01/00 16:31 adev Accepted_Events_Run_2489
  • 35136 03/01/00 16:16 adev Accepted_Events_Run_2488
  • 35134 03/01/00 16:08 adev Accepted_Events_Run_2487
  • 35133 03/01/00 16:05 F.Benmokhtar Number of events in Kin#28
  • 35129 03/01/00 15:07 bogdanw otr
  • 35128 03/01/00 14:57 bogdanw OTR
  • 35125 03/01/00 14:29 adev Accepted_Events_Run_2486
  • 35124 03/01/00 12:53 adev Accepted_Events_Run_2485
  • 35121 03/01/00 11:57 adev Accepted_Events_Run_2484
  • 35119 03/01/00 11:30 weinstei hrs magnet strip charts
  • 35118 03/01/00 11:28 weinstei beam strip chart
  • 35117 03/01/00 10:49 F.Benmokhtar Final rate for Kinematic 28
  • 35115 03/01/00 10:14 Garibaldi main control
  • 35113 03/01/00 09:20 F.Benmokhtar Correction of the Header file
  • 35112 03/01/00 09:06 adev Accepted_Events_Run_2482
  • 35108 03/01/00 08:59 F.Garibaldi e,e'p in kin. # 28
  • 35107 03/01/00 08:52 F. Garibaldi shift activity
  • 35105 03/01/00 08:01 M.Iodice Count estimates kin 28
  • 35103 03/01/00 07:44 M.Iodice DAQ problem
  • 35101 03/01/00 06:46 M.Iodice OTR measurement
  • 35099 03/01/00 06:22 adev Accepted_Events_Run_2480
  • 35097 03/01/00 06:10 M.Iodice gas system ALARM
  • 35096 03/01/00 06:07 C.Furget New Emiss for kin 28
  • 35094 03/01/00 05:43 adev Accepted_Events_Run_2479
  • 35092 03/01/00 04:54 C. Furget Counting estimate for kin 28
  • 35091 03/01/00 04:43 adev Accepted_Events_Run_2478
  • 35088 03/01/00 04:34 C. Furget Emiss for kin 28
  • 35087 03/01/00 04:24 M.Iodice hadron magnets strip chart
  • 35086 03/01/00 04:17 M.Iodice beam strip chart
  • 35085 03/01/00 04:16 M.Iodice Magnet strip charts
  • 35084 03/01/00 03:18 adev Accepted_Events_Run_2477
  • 35080 03/01/00 01:28 mitchell ep finished
  • 35079 03/01/00 00:35 mitchell start ep
  • 35078 03/01/00 00:01 x. jiang trigger file 1.28 gev downloaded.